OpenStudy (anonymous):

Determine the values of the trigonometric functions for the following. 1. t=4pi/3 2. t=3pi/4

OpenStudy (perl):

I assume these are angles

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

@perl do you know what csc, sec, and cot are?

OpenStudy (anonymous):

I know that sin=sq root-3/2, cos=-1/2, and tan=sq root-3, but I'm not sure how to get csc, sec, and cot